OpenStudy (anonymous):

three time a number s plus 4 times a number y is 1 more than 6 times the number s . solve for s

OpenStudy (anonymous):

3s + 4y - 1 = 6s 4y - 1 = 6s - 3s 4y - 1 = 3s 4y -1 ----- = s 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I have one more question

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

five times a number k minus 9 is two-third of a number j. solve for j

OpenStudy (anonymous):

what did you substract -1 instead of adding 1 on problem 1

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

you there, harkirat?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

first I will attend to the previous question three time a number s = 3s 4 times a number y = 4y so three time a number s plus 4 times a number y = 3s + 4y then 6 times the number s = 6s Your problem states "three time a number s plus 4 times a number y is 1 more than 6 times the number s" this means that 3s + 4y is one more than 6s this means that 6s is 1 less than 3s + 4y so if we subtract 1 from 3s + 4y then it becomes equal to 6s that is why 3s + 4y - 1 = 6s Understood ????

OpenStudy (anonymous):

Have patience, I am working on yr question only.....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Now coming to the second question five times a number k minus 9 is two-third of a number j. solve for j five times a number k minus 9 = 5k - 9 2 two-third of a number j = --- j 3 2 5k - 9 = = --- j 3 Multiplying both sides by 3 3(5k - 9) = 2j 15k - 27 = 2j 15k - 27 -------- = j 2

OpenStudy (anonymous):

Got a medal for me ................☺

OpenStudy (anonymous):

why muliplty by 3?

OpenStudy (anonymous):

2/3m+a=a+c solve for m

OpenStudy (anonymous):

2/3m+a=a+c solve for m

OpenStudy (anonymous):

2/3m+a=a+c solve for m

OpenStudy (anonymous):

I'll get you a medal

OpenStudy (anonymous):

to get rid of the 3 in the denominator on right hand side .....

OpenStudy (anonymous):

2/3m+a=a+c solve for m

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

why don't u give it a try.......... u will not learn unless u try to do them yrself....

OpenStudy (anonymous):

m=5/3c

OpenStudy (anonymous):

I did them, now I am checking them

OpenStudy (anonymous):

is m=5/3c correct?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

use the draw box to show me what u did...

OpenStudy (anonymous):

2/3m+a=a+c solve for m

OpenStudy (anonymous):

substract a then a will cancel out. mulipty by 3/2 to cancel out

OpenStudy (anonymous):

m=5/3c

OpenStudy (anonymous):

can you show me the answer?

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

see my draw box coming up....

OpenStudy (anonymous):